# Dynamic Financial Analysis (DFA) of General Insurers under Climate Change

### v1

@@ 2.4.1 Insurance costs
where $\mu^{\text{NC}}$ is the location parameter and $\phi$ is the dispersion parameter, both calibrated using historical data. Using a Tweedie distribution implies that claim frequency follows a Poisson distribution, while claim severity follows a Gamma distribution, reflecting the typically high-frequency, low-severity nature of non-catastrophe losses. Some studies have investigated the influence of weather on non-catastrophe claims (e.g., McGuire, 2008; Haug et al., 2011; Scheel et al., 2013; Reig Torra et al., 2023), but these typically require high-resolution daily or monthly municipal-level data, exceeding the usual granularity of DFA. Moreover, the long-term effect of weather on non-catastrophe claims is uncertain. For instance, McGuire (2008) found that same-day precipitation increases motor claims frequency, whereas lagged precipitation decreases it possibly due to cleaner road conditions (Eisenberg, 2004; McGuire, 2008). Consequently, we do not explicitly incorporate climate impacts in our modelling of non-catastrophe losses.
@@
$\tilde{X}_{t,(j)}^{\text{NC}}=X_{t}^{\text{NC}}\cdot\omega_{t}\cdot w_{j}\cdot\frac{\text{CPI}t}{\text{CPI}{s}},$ (2.17)
@@
The exposure growth modelling presented here does not consider the potential loss of business volume due to premium increases driven by climate change. This simplification relies on the assumption that household income growth will generally keep pace with rising premiums. While this may be less concerning under scenarios such as SSP 8.5—where both economic growth and climate risk are high—or SSP 2.6—where climate risk is low—the issue of affordability may become more significant under scenarios with weak income growth but elevated climate risk (e.g., SSP 7.0).
@@ 2.4.3 Reinsurance premiums
Reinsurers typically cover the extreme tail of insurers’ risk portfolios through excess-of-loss coverage, making them particularly vulnerable to large natural catastrophes. Such events can strain reinsurance capital and trigger hard markets with higher premiums, which is a trend expected to intensify as climate change increases catastrophe losses (Tesselaar et al., 2020). Capital constraint theory offers a common explanation, suggesting firms prefer to accumulate surplus through higher premiums rather than raise costly external capital, contributing to the persistence of hard market conditions (Winter, 1988, 1994; Dicks and Garven, 2022).
@@
We acknowledge that the reinsurance model presented here is a simplified representation, as it assumes a single reinsurer exclusively covering catastrophe losses in Australia. Despite its highly stylised nature, the model generally captures the inverse relationship between reinsurance capital and premium pricing dynamics. The impacts of climate change on the reinsurance cycle itself could be a separately interesting topic, and we will leave it for future research.
@@ 3.1.1 Data sources
In addition to the data sources mentioned above, we draw on Australian insurance market statistics from the General Insurance Performance Statistics database (APRA, 2024b) and the General Insurance Institution-Level Statistics database (APRA, 2024a). These data inform our market assumptions and support the calibration of the non-catastrophe loss model (see Section 2.4.1). Because no direct non-catastrophe loss data are publicly available, we derive industry-level non-catastrophe losses by subtracting the ICA-recorded catastrophe losses from the total industry losses reported in the General Insurance Performance Statistics database (APRA, 2024a), after joining those two datasets.
@@ 3.1.2 Calibration results
Table 3.1 presents the calibrated parameters for the selected model of each hazard type [^3] Each coefficient $\beta$ links a climate variable (superscript) to a distribution parameter (subscript). For example, $\beta_{(\lambda)}^{\text{rx5day}}$ denotes the sensitivity of frequency ($\lambda$) to rx5day, while $\beta_{(\mu)}^{\text{rx5day}}$ refers to its effect on the severity location parameter.. The full set of candidate models is detailed in Online Appendix A.2. Final models were chosen based on both physical relevance (see Section 2.2.2) and statistical performance. For example, in Table I, all flood frequency models show positive coefficients for precipitation, consistent with known flood mechanisms. Model 3, using rx5day as an extreme precipitation proxy, performed best by AIC and BIC and its covariate is statistically significant at the $5\%$ level. Thus, Model 3 is selected. As another case, although the atmospheric temperature coefficient is not statistically significant at the $5\%$ or $10\%$ levels, we retain it in the hailstorm model due to the theoretical importance of temperature effects (see Section 2.2.2).
@@
Table 3.3 presents the calibrated equity model parameters. $\alpha_{1}$ (sensitivity of operating profit growth to consumption growth) and $\sigma_{O}$ (its standard deviation) are calibrated using Australian operating profit data (Australian Bureau of Statistics, 2024a) and consumption growth data (World Bank Group, 2024). $\beta_{1}$ (sensitivity of excess equity returns to operating profit growth) and $\sigma_{X}$ (standard deviation of excess equity returns) are calibrated using operating profit data (Australian Bureau of Statistics, 2024a) and All-Ordinaries total returns (FactSet, 2024). Results indicate positive relationships between consumption and profit growth, and between profit growth and equity returns, consistent with expectations as discussed in Online Appendix B.
@@
Table 3.4 presents the market assumptions underlying our simulations, derived from general insurance market statistics (APRA, 2024c) and financial disclosures from individual insurers. In the projections, both excess and limit levels are adjusted for changes in GDP and CPI across future periods and under different climate scenarios.
@@
Table 3.5 summarises the additional assumptions used in the simulations. The target capital ratio is determined as the general insurance industry’s average ratio of eligible equity to the Minimum Capital Requirement (MCR), based on data from the APRA General Insurance Institution-Level Statistics database (APRA, 2024a). The uninsured-to-insured loss ratio is used to scale catastrophe insurance losses from the hazard module to uninsured economic damage, serving as an input to the equity model. This ratio is calculated as the average proportion of uninsured to insured losses over the historical period from 1985 to 2023, using data from the EM-DAT database (EM–DAT, 2023), since the ICA dataset records only insurance losses.
@@
$w_{rf}$ Allocation to risk-free assets $60\%$  [^5] The assumption is that the majority of bonds are invested in risk-free government securities, as insurers typically prefer government bonds over corporate bonds (OECD, 2023). Additionally, cash and deposits are considered nearly risk-free assets OECD (2023)
@@ 3.2.1 Climate and hazards
Based on the simulated climate variables, catastrophe losses for each hazard type are generated using the calibrated relationships between climate variables and hazard frequency and severity as described in Section 3.1.2. Overall, the projected normalised losses exhibit an increasing trend for most hazards, particularly under high-emission scenarios, with notable increases in both the mean and the upper tails of the loss distributions. Moreover, the results reveal considerable uncertainty, driven by both the variability in simulation of climate variables as shown in Appendix A and the heavy-tailed nature of catastrophe losses. The rising mean and volatility are expected to place upward pressure on both insurance and reinsurance premiums, as well as shocks to capital over time.
@@
Despite the simplified structure of our hazard models, the simulation results generally align with physical risk trends reported in the literature. Projected increases in bushfire and flood losses are consistent with previous findings (Chapter 12 of IPCC, 2021b, pp. 1805–1811), reflecting intensified extreme fire weather and increased extreme precipitation across much of Australia.
@@ 3.2.2 Investment returns
The simulation results for the cumulative investment returns, presented on a logarithmic scale, are shown in Figure 2(a)  [^6] Here, the log-scale cumulative investment return is defined as: $r^{\text{cum}}_{t}=\log\Bigl{(}\prod_{s=1}^{t}\bigl{(}1+r_{s}\bigr{)}\Bigr{)}$. This can also be written as: $r^{\text{cum}}_{t}=\log\Bigl{(}\prod_{s=1}^{t}\bigl{(}1+r_{s}\bigr{)}\Bigr{)}=\log\Bigl{(}\frac{V_{0}\prod_{s=1}^{t}(1+r_{s})}{V_{0}}\Bigr{)}=\log\Bigl{(}\frac{V_{t}}{V_{0}}\Bigr{)}$, where $V_{t}$ is the total investment value at time $t$.. The uncertainty bounds in the figure account for historical fluctuations in interest rates and equity returns, as well as the variability in climate-related damage to production available for consumption, as illustrated in Figure 3(a) and derived via (2.13).
The observed trends in cumulative investment returns are primarily driven by the economic growth assumptions underlying each SSP scenario (Figure 2(b)). Under SSP 8.5, the highest cumulative investment return is projected, driven by its robust economic growth assumption (O Neill et al., 2017), followed by SSP 2.6. Conversely, SSP 7.0 exhibits the weakest cumulative investment return, influenced by both its slow economic growth assumptions and high catastrophe damages, in line with its narrative (O Neill et al., 2017). The relatively high investment returns under SSP 8.5 and SSP 2.6 are likely to accelerate capital accumulation. By contrast, the lower returns, coupled with high catastrophe losses in SSP 7.0 (see Figures 3(b) and 3.1), are expected to slow the pace of capital accumulation under this scenario.
@@
The projected damage ratios from our climate-dependent DFA model are also compared with estimates from existing literature; a detailed discussion can be found in Online Appendix C.
Figure 3.2: Simulated (log) compounded investment returns (a) and projected compounded real GDP growth in Australia (b). Panel (a) shows the simulated compounded investment returns generated from the economic growth assumptions and the simulated hazard losses underlying each scenario. Panel (b) presents compounded real GDP growth projections derived from the SSP database (Riahi et al., 2017).
@@ 3.2.3 Premiums and underwriting losses
The normalized gross premiums associated with catastrophe (CAT) coverage are presented in Figure 4(a). An overall increasing trend is found, driven by the rising risk of most hazards as shown in Figure 3.1. Gross CAT premiums reach their highest levels under SSP 8.5, particularly in the later projection period, followed by SSP 7.0, SSP 4.5, and SSP 2.6, in line with the physical risk narratives underlying those scenarios as described in Section 2.1. In addition to this general upward trend, significant inter-annual variability is observed, reflecting the internal climate variability of the underlying climate variables as shown in AppendixA.
Moreover, the proportion of CAT premiums relative to total general insurance premiums is expected to increase relative to the historical levels, with a more pronounced increase under high-emission scenarios. This is illustrated in Figure 4(b), which compares projected CAT premium proportions with historically observed CAT loss proportions (used here as a proxy for historical CAT premium proportions) derived from the General Insurance Performance Database (APRA, 2024c) and the ICA dataset. These findings suggest that catastrophe losses will have an increasingly significant impact on the underwriting performance of general insurers under the influence of climate change. Consequently, managing CAT exposures will become a more critical component of portfolio management for general insurers.
@@
Figure 3.4: Gross premiums (normalised) associated with catastrophe cover, shown as dollar values (a) and as a proportion of total gross premiums (b). Panel (a) presents the normalised values of calculated gross catastrophe (CAT) premiums based on (2.19). Panel (b) shows the projected CAT premiums as a proportion of total premiums, compared with historically observed CAT losses as a share of total insurance losses, based on data from the APRA database and the ICA dataset. Both panels indicate an increasing trend in CAT premiums—both in dollar terms and as a share of total premiums—particularly under high-emission scenarios.

### v2

@@ 2.4.1 Insurance costs
where $\mu^{\text{NC}}$ is the location parameter and $\phi$ is the dispersion parameter, both calibrated using historical data. Using a Tweedie distribution implies that claim frequency follows a Poisson distribution, while claim severity follows a Gamma distribution, reflecting the typically high-frequency, low-severity nature of non-catastrophe losses. Some studies have investigated the influence of weather on non-catastrophe claims (e.g., McGuire, 2008; Haug et al., 2011; Scheel et al., 2013; Reig Torra et al., 2023), but these typically require high-resolution daily or monthly municipal-level data, exceeding the usual granularity of DFA. Moreover, the long-term effect of weather on non-catastrophe claims is uncertain. For instance, McGuire (2008) found that same-day precipitation increases motor claims frequency, whereas lagged precipitation decreases it possibly due to cleaner road conditions (Eisenberg, 2004; McGuire, 2008). Consequently, we do not explicitly incorporate climate impacts in our modelling of non-catastrophe losses, and the integration of climate covariates is left for future research as more data become available.
@@
$\tilde{X}_{t,(j)}^{\text{NC}}=X_{t}^{\text{NC}}\cdot\omega_{t}\cdot w_{j}\cdot\frac{\text{CPI}_{t}}{\text{CPI}_{s}},$ (2.17)
@@
The exposure growth modelling presented here assumes that business volume grows in line with population growth and does not consider the potential loss of business volume due to premium increases driven by climate change. This simplification relies on the assumption that household income growth will generally keep pace with rising premiums. While this may be less concerning under scenarios such as SSP 8.5—where both economic growth and climate risk are high—or SSP 2.6—where climate risk is low—the issue of affordability may become more significant under scenarios with weak income growth but elevated climate risk (e.g., SSP 7.0). As a potential extension, one could incorporate the influence of GDP-per-capita growth relative to premium growth on premium affordability, and its implications on insurance demand.
@@ 2.4.3 Reinsurance premiums
Reinsurers typically cover the extreme tail of insurers’ risk portfolios through excess-of-loss coverage, making them particularly vulnerable to large natural catastrophes. Such events can strain reinsurance capital and trigger hard markets with higher premiums, which is a trend expected to intensify as climate change increases catastrophe losses (Tesselaar et al., 2020). Capital constraint theory offers a common explanation, suggesting firms prefer to accumulate surplus through higher premiums rather than raise costly external capital (Winter, 1988, 1994; Dicks and Garven, 2022).
@@
In modelling insurance premiums, we assume that insurance contracts are repriced annually to reflect changes in risk. While this assumption is appropriate for the scope of this paper, which is to provide a baseline framework, it should be noted that, in practice, premium adjustments may lag behind changes in risk due to underwriting cycles (see, e.g., Winter, 1994), market competition (see, e.g., Taylor, 2008), and rate regulation (see, e.g., Weiss et al., 2010; Taylor, 2008). Such pricing delays may further increase insolvency risk through periods of underpricing. Future research could incorporate these factors into the modelling of premium dynamics and assess their implications for the risk–return profile of general insurers under various SSP scenarios, relative to the baseline results presented in this paper.
We acknowledge that the reinsurance model presented here is a simplified representation, as it assumes a single reinsurer exclusively covering catastrophe losses in Australia and does not allow reinsurance prices to fall below their risk-based premium levels. In practice, the reinsurance market typically involves multiple participants, with reinsurers operating globally (BeInsure Media, 2025), offering both catastrophe and non-catastrophe coverage, and prices that can occasionally fall below their risk-based premium levels (Enz, 2002; Meier and Outreville, 2010). Despite its highly stylised nature, the model generally reflects the prudence that reinsurance prices are likely to rise above risk-based premiums as reinsurance capital tightens. However, it should not be interpreted as a comprehensive modelling of reinsurance market cycles due to its simplified assumption mentioned above. The potential impacts of climate change on the reinsurance cycle itself constitute an interesting topic for future research.
@@ 3.1.1 Data sources
In addition to the data sources mentioned above, we draw on Australian insurance market statistics from the General Insurance Performance Statistics database (APRA, 2024c) and the General Insurance Institution-Level Statistics database (APRA, 2024a). These data inform our market assumptions and support the calibration of the non-catastrophe loss model (see Section 2.4.1). Because no direct non-catastrophe loss data are publicly available, we derive industry-level non-catastrophe losses by subtracting the ICA-recorded catastrophe losses from the total industry losses reported in the General Insurance Performance Statistics database (APRA, 2024a), after joining those two datasets.
@@ 3.1.2 Calibration results
Table 3.1 presents the calibrated parameters for the selected model of each hazard type [^4] Each coefficient $\beta$ links a climate variable (superscript) to a distribution parameter (subscript). For example, $\beta_{(\lambda)}^{\text{rx5day}}$ denotes the sensitivity of frequency ($\lambda$) to rx5day, while $\beta_{(\mu)}^{\text{rx5day}}$ refers to its effect on the severity location parameter.. The full set of candidate models is detailed in Online Appendix A.3. Final models were chosen based on both physical relevance (see Section 2.2.2) and statistical performance. For example, in Table J, all flood frequency models show positive coefficients for precipitation, consistent with known flood mechanisms. Model 3, using rx5day as an extreme precipitation proxy, performed best by AIC and BIC and its covariate is statistically significant at the $5\%$ level. Thus, Model 3 is selected. As another case, although the atmospheric temperature coefficient is not statistically significant at the $5\%$ or $10\%$ levels, we retain it in the hailstorm model due to the theoretical importance of temperature effects (see Section 2.2.2).
@@
Table 3.3 presents the calibrated equity model parameters. $\alpha_{1}$ (sensitivity of operating profit growth to consumption growth) and $\sigma_{O}$ (its standard deviation) are calibrated using Australian operating profit data (Australian Bureau of Statistics, 2024a) and consumption growth data (World Bank Group, 2024). $\beta_{1}$ (sensitivity of excess equity returns to operating profit growth) and $\sigma_{X}$ (standard deviation of excess equity returns) are calibrated using operating profit data (Australian Bureau of Statistics, 2024a) and All-Ordinaries total returns (FactSet, 2024). Results indicate positive relationships between consumption and profit growth, and between profit growth and equity returns, consistent with expectations as discussed in Online Appendix C.1.
@@
Table 3.4 presents the market assumptions underlying our simulations, derived from general insurance market statistics (APRA, 2024b) and financial disclosures from individual insurers. In the projections, both excess and limit levels are adjusted for changes in GDP and CPI across future periods and under different climate scenarios.
@@
Table 3.5 summarises the additional assumptions used in the simulations. The target capital ratio is determined as the general insurance industry’s average ratio of eligible equity to the Minimum Capital Requirement (MCR), based on data from the APRA General Insurance Institution-Level Statistics database (APRA, 2024a). The uninsured-to-insured loss ratio is used to scale catastrophe insurance losses from the hazard module to uninsured economic damage, serving as an input to the equity model. This ratio is calculated as the average proportion of uninsured to insured losses over the historical period from 1985 to 2023, using data from the EM-DAT database (EM–DAT, 2023), since the ICA dataset records only insurance losses. It should also be noted that this ratio represents only the proportion of uninsured to insured damages to assets, which are primarily properties and motor vehicles that tend to have relatively high insurance coverage due to regulatory and mortgage requirements (Melser et al., 2024), during natural disasters. It does not, however, reflect the overall insurance penetration of the broader economy. Moreover, this ratio may vary across countries and can be affected by differences in the quality and reporting protocols of hazard impact data across regions (EM-DAT, 2023).
@@
$w_{rf}$ Allocation to risk-free assets $0\% 66$6The assumption is that the majority of bonds are invested in risk-free government securities, as insurers typically prefer government bonds over corporate bonds (OECD, 2023). Additionally, cash and deposits are considered nearly risk-free assets OECD (2023)
@@ 3.2.1 Climate and hazards
Based on the simulated climate variables, catastrophe losses for each hazard type are generated using the calibrated relationships between climate variables and hazard frequency and severity as described in Section 3.1.2. Overall, the projected normalised losses exhibit an increasing trend for most hazards, particularly under high-emission scenarios, with notable increases in both the mean and the upper tails of the loss distributions. In addition, the projected losses exhibit substantial interannual variability. For example, a pronounced peak in simulated flood losses is observed around 2047, particularly at the $95^{\text{th}}$ percentile. This peak is primarily driven by the corresponding increase in extreme precipitation (see Online Appendix B.1), which may be explained by the projected rise in the frequency of extreme ENSO events under high-emission scenarios reported in the literature (see, e.g., Cai et al., 2015; Heede and Fedorov, 2023). Moreover, the results reveal considerable uncertainty, driven by both the variability in simulation of climate variables as shown in Appendix A and the heavy-tailed nature of catastrophe losses. The rising mean and volatility are expected to place upward pressure on both insurance and reinsurance premiums, as well as shocks to capital over time.
@@
Despite the simplified structure of our hazard models, the simulation results generally align with physical risk trends reported in the literature. Projected increases in bushfire (with the upward trend more clearly illustrated in the extended simulation results presented in Online Appendix B.2) and flood losses are consistent with previous findings (Chapter 12 of IPCC, 2021b, pp. 1805–1811), reflecting intensified extreme fire weather and increased extreme precipitation across much of Australia.
@@ 3.2.2 Investment returns
The simulation results for the cumulative returns on the total investment portfolios, presented on a logarithmic scale, are shown in Figure 2(a)  [^7] Here, the log-scale cumulative investment return is defined as: $r^{\text{cum}}_{t}=\log\Bigl(\prod_{s=1}^{t}\bigl(1+r_{s}\bigr)\Bigr)$. This can also be written as: $r^{\text{cum}}_{t}=\log\Bigl(\prod_{s=1}^{t}\bigl(1+r_{s}\bigr)\Bigr)=\log\Bigl(\frac{V_{0}\prod_{s=1}^{t}(1+r_{s})}{V_{0}}\Bigr)=\log\Bigl(\frac{V_{t}}{V_{0}}\Bigr)$, where $V_{t}$ is the total investment value at time $t$.. The uncertainty bounds in the figure account for historical fluctuations in interest rates and equity returns, as well as the variability in climate-related damage to production available for consumption, as illustrated in Figure 3(a) and derived via (2.12).
The observed trends in cumulative investment returns are primarily driven by the economic growth assumptions underlying each SSP scenario (Figure 2(b)). Under SSP 8.5, the highest cumulative investment return is projected, driven by its robust economic growth assumption (O Neill et al., 2017), followed by SSP 2.6. Conversely, SSP 7.0 exhibits the weakest cumulative investment return, influenced by both its slow economic growth assumptions and high catastrophe damages, in line with its narrative (O Neill et al., 2017). The relatively high investment returns under SSP 8.5 and SSP 2.6 are likely to accelerate capital accumulation. By contrast, the lower returns, coupled with high catastrophe losses in SSP 7.0 (see Figures 3(b) and 3.1), are expected to slow the pace of capital accumulation under this scenario. A sensitivity analysis using alternative economic damage assumptions has also been conducted (see Online Appendix G), indicates broadly consistent patterns with those reported in this section.
A breakdown analysis of investment returns for the general and brown portfolios is provided in Online Appendix C.2. Overall, the transition stress remains relatively mild, as the SSP scenarios considered do not assume abrupt or disorderly transitions, as noted in Remark 2.1.
@@
The projected damage ratios from our climate-dependent DFA model are also compared with estimates from existing literature; a detailed discussion can be found in Online Appendix D and G.
Figure 3.2: Simulated (log) compounded investment returns (a) and projected compounded real GDP growth in Australia (b). Panel (a) shows the simulated compounded returns on the total investment portfolios generated from the economic growth assumptions and the simulated hazard losses underlying each scenario. Panel (b) presents compounded real GDP growth projections derived from the SSP database (Riahi et al., 2017).
@@ 3.2.3 Premiums and underwriting losses
The normalized gross premiums associated with catastrophe (CAT) coverage are presented in Figure 4(a). An overall increasing trend is found, driven by the rising risk of most hazards as shown in Figure 3.1. Gross CAT premiums reach their highest levels under SSP 8.5, particularly in the later projection period, followed by SSP 7.0, SSP 4.5, and SSP 2.6, in line with the physical risk narratives underlying those scenarios as described in Section 2.1. The increase in premiums may also create affordability issues and, as noted in Remark 2.8, affect insurance demand. This effect is noted as a limitation of the current paper but is encouraged for future research. In addition to this general upward trend, significant inter-annual variability is observed, reflecting the internal climate variability of the underlying climate variables as shown in AppendixA.
Moreover, the proportion of CAT premiums relative to total general insurance premiums is expected to increase relative to the historical levels, with a more pronounced increase under high-emission scenarios. This is illustrated in Figure 4(b), which compares projected CAT premium proportions with historically observed CAT loss proportions (used here as a proxy for historical CAT premium proportions) derived from the General Insurance Performance Database (APRA, 2024b) and the ICA dataset. These findings suggest that catastrophe losses will have an increasingly significant impact on the underwriting performance of general insurers under the influence of climate change. Consequently, managing CAT exposures will become a more critical component of portfolio management for general insurers.
@@
The results also suggest that the percentage growth in reinsurance premiums exceeds that of primary insurance, reflecting both rising physical risks and adjustments to tightening reinsurance capital. The modelled uplift relative to the base level is considered as conservative due to the simplified assumptions outlined in Remark 2.9. Nevertheless, it is assumed that primary insurers will continue to purchase reinsurance despite higher prices; however, rising premiums may dampen demand, although this effect could be partly offset by the strengthening risk-transfer demand as climate risk intensifies (Lloyd’s, 2025). The dynamics of these effects are left for future research.
Figure 3.4: Gross premiums (normalised) associated with catastrophe cover, shown as dollar values (a) and as a proportion of total gross premiums (b). Panel (a) presents the normalised values of calculated gross catastrophe (CAT) premiums based on (2.19). Panel (b) shows the projected CAT premiums as a proportion of total premiums, compared with historically observed CAT losses as a share of total insurance losses, based on data from the APRA database and the ICA dataset. Both panels indicate an increasing trend in CAT premiums–both in dollar terms and as a share of total premiums–particularly under high-emission scenarios.